Two springs are

joined and connected to a block of

mass 0.245 kg that is set oscillating

over a frictionless floor. The springs

each have spring constant k $

6430 N/m. What is the frequency of

the oscillations?


1
Expert's answer
2021-12-28T13:36:32-0500

Let's suppose that the two springs connected in series with each other. Then, we can find its total spring constant as follows:


"\dfrac{1}{k_{tot}}=\dfrac{1}{k_1}+\dfrac{1}{k_2},""\dfrac{1}{k_{tot}}=\dfrac{1}{k}+\dfrac{1}{k}=\dfrac{2}{k},""k_{tot}=\dfrac{k}{2}."

Finally, we can find the frequency of oscillations as follows:


"f=\dfrac{1}{2\pi}\sqrt{\dfrac{k_{tot}}{m}}=\dfrac{1}{2\pi}\sqrt{\dfrac{k}{2m}},""f=\dfrac{1}{2\pi}\sqrt{\dfrac{6430\ \dfrac{N}{m}}{2\times0.245\ kg}}=18.2\ Hz."